40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 218 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  5924   Thu Nov 17 11:51:14 2011 JenneUpdateEnvironmentIncandescent vs. fluorescent lights?

I'm just on an elog roll this morning...

Again while poking around inside the IFO room, I noticed that they have replaced all of our incandescent lights with CFLs.  Do we care?  The point of having the incandescent lights on a separate switch from the big fluorescent lights was so that we could have only 60Hz lines, not wide-band noise if we want the lights on while locking. 

I'm not sure that we actually care, because more often we just turn off all the lights while trying to do serious locking, but if we do care, then someone needs to ask the custodial staff (or someone else?) to undo the change.

  5923   Thu Nov 17 11:35:27 2011 KojiUpdateRF SystemStochmon?

The  Stochmon channels for 11&55MHz have been reasonably working since last night.

The output is not yet calibrated as the RF power detector has a strange scaling.
I am analyzing the calibration data.

  5922   Thu Nov 17 11:27:58 2011 JenneUpdatePSLHEPA turned down

I was measuring things to see how big my adapter plate needs to be, and I decided that we'd had enough days of the HEPA being on full blast, so I turned it down to 50, from 100.  I think it's been on full since Katrin was working on the Y-green beat a week or so ago.

  5921   Thu Nov 17 11:04:02 2011 JenneUpdateRF SystemStochmon?

Is there an update on Stochmon?  Are the signals acquired somewhere already?  What's the current deal-io?  The new EOM mount should be here later today, and I'm jazzed to start checking how my EOM box helps (hopefully) the amount of RFAM we see. 

I'll start making the adapter plate while I wait...

  5920   Thu Nov 17 03:46:52 2011 kiwamuUpdateGreen LockingPSL doubling had been diabled

I found that the temperature controller of the PSL doubling oven had been disabled.

Because of that I took a little bit long time to recover the beat-note.
I have no idea why its been disabled.
I turned it on to make the PSL green beam bright,
Also the I-parameter of the PID temperature control was too big
and because of that a big overshoot in the temperature happened (overshoot of ~ 5 deg !).
So I decreased the I-parameter from 175 to 85 (250 is the maximum).
Now the intensity of the green light seems reasonably bright and stable.
  5919   Wed Nov 16 23:50:40 2011 DenUpdateAdaptive Filteringseismic noise injection

[Micro, Den]

Analyzing coherence of seismic noise and mode cleaner length we've figured out that at some days the coherence below 1 Hz is still present. For example, at Nov 13 we can see some coherence compared to most other dates when we are not able to see coherence as shown on the figure. On the top plot - psd of MC_L and GUR1_X at Nov 13 (red and blue) and Nov 15 (black and cyan). On the bottom plot is presented coherence between MC_L and GUR1_X on Nov 13 (red) and Nov 15 (black)

datespsd.jpg

datescoh.jpg

We can divide the psd plot for 2 parts - below 1 Hz and above 1 Hz. Above 1 Hz seismic noise on Nov 15 (cyan) was higher then on Nov 13 (blue) and correspondently MC_L at that region was higher on Nov 15. Below 1 Hz seismic noise was higher on Nov 13 but MC_L is still lower that on Nov 15. That is surprising. From the coherence plot we can say that once we have some more seismic noise than usually, we immediately see coherence.

Because of this we wanted to find out the level of the X noise that makes seismic noise invisible. We injected seismic noise by doing smooth physical exercises near MC_2 (1.5 m and 3 m apart). The MC_2 was in lock during the experiment.

injectionpsd.jpg

injectioncoh.jpg

In the coherence plot we can see that coherence between GUR1_X and MC_L increased with noise injection. The highest coherenced we obtained sittind down and standing up smoothly near MC_2 at distance 1.5 m. We did not want to come clother and break the lock. This measurement tells us that the X noise is approximately 3-4 times higher than seismic noise in the range 0.1 - 1 Hz. That means that it is approximately 1e-6 - 1e-8 m/sqrt(Hz) in this region. This noise goes down at frequencies from 2 Hz and not seen because of seismic noise. Actually, seismic noise can be filtered out with the Wiener filter and then we'll see the spectrum of X noise.

We now try to figure out the method to estimate the contribution of OSEM noise to the X noise.

  5918   Wed Nov 16 21:01:08 2011 JenneUpdateTreasureeom box

I made a super sweet new foam box for our EOM.  It's awesome, and should be reasonably easy to duplicate.  Check out the PHOTOS!

Notes:

* I didn't think I was going to cover the inside of the box at first, since the foam is non-fuzz-generating, but Koji suggested it would be a good idea anyway.  The foam was cut perfectly to the EOM, so adding the tape inside makes it a tight fit.  Especially height-wise...leave a little space next time.

* To cover the insides of the optical path holes, do it in 2 parts.  One half-cylinder, and then another.  Way easier than trying to do the whole thing at once.  Also, pre-cut the tabs on both sides of the foil before inserting.  Then you just have to grab the tabs with tweezers and flatten them, and they hold the aluminum tape in place. 

* Having 1" wide, 2" wide and 3" wide aluminum tape was handy.  3" to make the top, 2" for the sides, and 1" for the inside of the holes. 

  5917   Wed Nov 16 20:30:27 2011 not KojiUpdateIOOMC unlocked and misaligned.

Quote:

Actually, do we need to reset the filter history at every lock loss of the MC?

Those DC offsets were necessary to keep the alignment good just until the MC is unlocked.
So if we keep the history, we can maintain the good alignment.

 I suspect the integrators get fed a huge wrong signal on lockloss. Clearing the history on the trans DOFs when the MC was badly aligned gets it nicely aligned again. I switched off the alignment transmission DOFs for now.

  5916   Wed Nov 16 18:14:09 2011 KojiUpdateIOOMC unlocked and misaligned.

Actually, do we need to reset the filter history at every lock loss of the MC?

Those DC offsets were necessary to keep the alignment good just until the MC is unlocked.
So if we keep the history, we can maintain the good alignment.

  5915   Wed Nov 16 17:40:48 2011 MirkoUpdateIOOMC unlocked and misaligned.

MC fell out of lock and was then quite badly misaligned. Mostly in pitch. I realigned it and it locked ok.

Turns out the MC falls often out of lock when the WFS servo comes on. I think the MC2_Trans history is not cleared on lockloss. I cleared it manually and realigned. Seems fine for now.

  5914   Wed Nov 16 17:29:46 2011 kiwamuUpdateGreen LockingSome updates on the Y end green PDH
Quote from #5894

 (Things to be done)

   [DONE]   1.1 Measurement of the arm fluctuation => to allow re-designing the servo shape
   [DONE]   1.2 temporary SR560 servo
   [ONGOING]1.3 Sanity checks on the modulation depth, reflectivity, PD dark noise and etc.,
   [DONE]  1.4 Make the servo more robust
   [DONE]  1.5 Some modifications on the medm screens
   [NOTYET]   1.6 Activation of the temperature feedback through the realtime digital control

Some updates on the Y end green PDH lock

(Measurement of the Y arm fluctuation)

In order to design the PDH box's servo shape we wanted to measure the Y arm fluctuation.
Here is the spectrum taken by looking at the control signal before the laser PZT.
 
 Yarm_fluctuation.png
 The scale of the Y axis is calibrated by using the PZT response of 5 MHz/V.
Above 10 Hz the spectrum shows 1/f noise which I believe the laser frequency noise.
 

(Temporary servo setup)

 We have found that the servo shape was not enough (#5890) to well-suppress the fluctuation shown above.
 Since the Newfocus fast servo box only makes 1/f shape, the error signal wasn't suppressed within the linear range.
So I have added an SR560 in the other input of the Newfocus servo box to make the filter shape 1/f^2.
Then the lock became more solid and the reflected DC light in time series is now much flat if the alignments are good.
I will post the servo shape and diagram later.

(Sanity checks)

 I looked at the reflected DC light when the laser was kept locked.
The reflectivity of the Y arm cavity went down to about 30% and this is good because it is supposed to be 27.5% when it is locked according the spec.
This means the mode-matching is not so bad.
  5913   Wed Nov 16 17:03:19 2011 KojiUpdateIOOMC2 Shifted in Pitch, corrected by adjusting the pitch bias

MC was not locked for more than 5 hours because of the misalignment.

Noticed that MC2 WFS feedback filters had big outputs (particularly in Pitch).
They were reset to zero.

MC2 was aligned and recovered the lock. Once the WFS is engaged, the transmission returned to the uisual value.

  5912   Wed Nov 16 14:34:18 2011 steveUpdatePSLIOO beam moves in pitch

C1:IOO-QPD_ANG_VERT beam movement  in 1 degree C temp change in 3 hrs

 

Attachment 1: iooqpds.png
iooqpds.png
  5911   Wed Nov 16 12:21:33 2011 KojiUpdateeloggooglebot (Re: restarted)

- elogd itself is a sort of web server which has no freedom to put our own file, we can not put robots.txt

- If we include elog using proxy in the usual web tree of Apache, we can put robots.txt at the root.

- So far, if we prevent "page0" browse by google, we will be saved for a while.

- It seems that the indexing is set to be refused by the following meta tags. But it does not prohibit googlebot to use "page0" URL, of course.

<META NAME="ROBOTS" CONTENT="NOINDEX, NOFOLLOW">
  5910   Wed Nov 16 10:53:35 2011 SureshUpdateIOOMC2 Shifted in Pitch, corrected by adjusting the pitch bias

Quote:

[Steve, Suresh]

    Steve went over to the MC2 walkway and stepped over the barrier to pick up some stuff there.  MC2 stack shifted and MC2 pitch as off.  MC unlocked and could not relock till the MC2 pitch bias was readjusted

previous MC2PIT reading: 3.6235           current MC2PIT reading:  3.9565

Without the WFS the MC to PSL alignment is poor, but it is largely due to a shift in the MC and not a shift in the PSL beam.  We know this 'coz the shift in the DC spot positions on WFS (when the MC is unlocked) is not significant nor is the shift on the C1:IOO-QPD.  When WFS loops are engaged the MC optics are turned to optimise the PSL to MC alignment, but the shift is large at the moment.

(Sorry Mirko your measurement could not be completed.  The MC unlocked in the middle)

Please Note:  If you need to access the blocked off area near MC2 stack, do not step over the barrier.  The disturbance is too great and the MC2 stack will shift.  Instead please move the barrier aside and walk as gently as possible near it, taking care not to touch the MC2 Chamber.

 

Apparently the MC2 stack had not finished shifting.   The MC unlocked while Steve was working on the PSL table installing the mirror for IOO_QPD and then it could not relock.  So I moved the MC2 once again in Pitch.  The current status of the sliders is here

C1IOO_MC_ALIGN.png

 

Yesterday I fixed the yellow buttons on the MC_ALIGN and MCLOCK screens.  They use the new updatesnap script  .  Could we also add a couple of lines to this script so that eveytime we save a snap shot the various values are written(appended) to a text file?  That way we do not need to depend solely on the conlog, which is quite slow.

 

  5909   Wed Nov 16 10:25:57 2011 steveUpdateSUSPRM damping restored

The  PRM lost damping about a day ago. It was restored.

  5908   Wed Nov 16 10:13:13 2011 jamieUpdateelogrestarted

Quote:

Basically, elog hangs up by the visit of googlebot.

Googlebot repeatedly tries to obtain all (!) of entries by specifying  "page0" and "mode=full".
Elogd seems not to have the way to block an access from the specified IP.
We might be able to use http-proxy via apache. (c.f. http://midas.psi.ch/elog/adminguide.html#secure)

 There are much more simple ways to prevent page indexing by googlebot: http://www.google.com/support/webmasters/bin/answer.py?answer=156449

However, I really think that's a less-than-idea solution to get around the actual problem which is that the elog software is a total piece of crap.  If google does not index the log then it won't appear in google search results.

But if there is one url that when requested causes the elog to crash then maybe it's a better solution to cut of just that url.

  5907   Wed Nov 16 10:11:20 2011 steveUpdatePSLIOO angle & pos qpd centered

Quote:

This moring I centered the IOO Angle QPD. The IOO Pos QPD was not centered. The existing layout does not allow the beam centering of the Pos qpd without misaligning the MC

input. We have to add an aditional steering mirror. I will do that tomorrow morning.

 I added the steering mirror for Pos  and centered both qpds

Attachment 1: iooqpds.png
iooqpds.png
  5906   Wed Nov 16 10:08:17 2011 SureshUpdateIOOEffect of turning on the MC2_TRANS_PIT and YAW loops in ASC

I turned on the two remaining loops in the ASC system to see if we can lock.   I put in some ones into the WFS_OUTPUT matrix

WFS_OUTMATRIX.png

and locked the MC2_TRANS_PIT and MC2_TRANS_YAW loops.

The effect of doing so is visible in the error signals.  The black loops are with all ASC loops off, Blue traces are with the WFS1 and 2 loops locked and Red traces are with all loops locked.  I took the red traces to a lower frequency to see if the suppression of the error signals at low frequencies is disturbed by the switching on of the MC2_TRANS loops.  They seem to be working fine without adding any perturbation above the UGF.

WFS_servo_err_20111115.png

I measured the  Transfer Function coefs (at 10Hz using the WFS Lockins)  with MC2_TRANS loops locked in this rudimentary fashion

  WFS1P WFS2P MC2TP WFS1Y WFS2Y MC2TY
MC1P -23.8541 15.2501 -24.3470 -3.3166 -2.0473 -0.1202
MC2P 29.7402 54.7689 29.5102  -0.2922 -17.4226 0.0310
MC3P 34.3612 10.7279 33.9650 6.6582 -4.0892 0.2333
MC1Y 0.9510 -6.3929 0.8722 -98.2414 -82.9129 -4.2802
MC2Y 12.0673 6.1708 11.9502 237.1172 20.7970 14.6480
MC3Y -0.8498 2.8712 -1.4195 -20.6031 111.2531 -1.5234

 

The green and blue bits are the only relevant parts since we ignore the off diagonal parts.  And most of these off diagonal coefs are indeed quite small (<5% of the max).  I have marked the not-so-small ones in yellow.

I then calculated the output matrix elements in two different ways.

a) Using a null vector in the place of MC_DoF --> MC2_TRANS transfer coefs.  The output matrix we get is

 

  WFS1P WFS2P Null Vector
MC1P -1.0000 0.8271  -0.8880
MC2P 0.0962 1.0000  0.4431
MC3P 0.9306 -0.2913  -1.0000

 

  WFS1Y WFS2Y Null Vector
MC1Y -0.2340 -0.5840 1.0000
MC2Y 1.000o -0.1551  0.4714
MC3Y -0.3613 1.0000 0.6571

 

b) Without using the null vector.  i.e. using the MC_DoF --> MC2_TRANS transfer coefs and inverting the full matrix.  The output matrix we get is

 

   WFS1P WFS2P  MC2TP
 MC1P  0.1471  -0.8880  0.8655
 MC2P  1.0000  0.4431  -0.4369
 MC3P  -0.7634  -1.0000  1.0000

 

  WFS1Y WFS2Y MC2TP
MC1Y 0.1401 1.0000 -1.0000
MC2Y 0.1449 0.4714 -0.3627
MC3Y 1.0000 0.6571 -0.6775

 

I plan to try out these two output matrices and measure the OL TFs of the MC2_TRANS and see if we can include these into ASC in a useful fashion.

Attachment 1: WFS_OUTMATRIX.png
WFS_OUTMATRIX.png
  5905   Wed Nov 16 09:21:56 2011 SureshUpdateIOOMC2 Shifted in Pitch, corrected by adjusting the pitch bias

[Steve, Suresh]

    Steve went over to the MC2 walkway and stepped over the barrier to pick up some stuff there.  MC2 stack shifted and MC2 pitch as off.  MC unlocked and could not relock till the MC2 pitch bias was readjusted

previous MC2PIT reading: 3.6235           current MC2PIT reading:  3.9565

Without the WFS the MC to PSL alignment is poor, but it is largely due to a shift in the MC and not a shift in the PSL beam.  We know this 'coz the shift in the DC spot positions on WFS (when the MC is unlocked) is not significant nor is the shift on the C1:IOO-QPD.  When WFS loops are engaged the MC optics are turned to optimise the PSL to MC alignment, but the shift is large at the moment.

(Sorry Mirko your measurement could not be completed.  The MC unlocked in the middle)

Please Note:  If you need to access the blocked off area near MC2 stack, do not step over the barrier.  The disturbance is too great and the MC2 stack will shift.  Instead please move the barrier aside and walk as gently as possible near it, taking care not to touch the MC2 Chamber.

  5904   Wed Nov 16 08:57:08 2011 SureshUpdateIOOMC WFS Servo OLG data and fits

I measured the Transfer Functions between from IN2 to IN1 on the WFS1PIT, WFS2PIT, WFS1YAW and WFS2YAW servo loops. 

Then I used the foton filter profiles of the servo filters in the loop and added another one to simulate the pendulum to generate a reasonable fit to the data.  Only the pendulum filter was hand tweaked since the PIT and YAW pendula have different resonant frequencies.

The filter modules included are:

1) Integrator: zpk([0.8],[0],0.8,"n")

2) Phase lead: zpk([0.8],[100,100],1,"n")

3) 45 deg filter: zpk([1:10],[3,30],1,"n")

4) ELP28: ellip("LowPass",5,1,50,28)

5)Pendulum: zpk([ ],0.03+i*0.82;0.03+i*0.82;],1,"n"  (for YAW)

5)Pendulum: zpk([ ],0.05+i*0.68;0.05+i*0.68;],1,"n"  (for PIT)

The data and fits are below.   The UGF is around 2 to 3 Hz and there is no servo bump at this gain setting.  The fits are poor at and below the resonance because the coherence was poor at these frequencies.  I will have to do a swept sine measurement for these low frequencies.

WFS1PITservo.png  WFS2PITservo.pngWFS1YAWservo.png  WFS2YAWservo.png

  5903   Wed Nov 16 02:36:35 2011 KojiUpdateelogrestarted

Basically, elog hangs up by the visit of googlebot.

Googlebot repeatedly tries to obtain all (!) of entries by specifying  "page0" and "mode=full".
Elogd seems not to have the way to block an access from the specified IP.
We might be able to use http-proxy via apache. (c.f. http://midas.psi.ch/elog/adminguide.html#secure)

Or can we tweak the source of elod such that it does not accept "page0"?


GET /40m/page0?mode=full&new_entries=1 HTTP/1.1
Host: 131.215.115.52:8080
Connection: Keep-alive
Accept: */*
From: googlebot(at)googlebot.com
User-Agent: Mozilla/5.0 (compatible; Googlebot/2.1; +http://www.google.com/bot.html)
Accept-Encoding: gzip,deflate

 

Quote:

Elog was hanging. Restarted it with script.

 

  5902   Wed Nov 16 01:45:37 2011 ZachUpdateelogrestarted

Elog was hanging. Restarted it with script.

  5901   Tue Nov 15 23:44:44 2011 MirkoUpdateCDSC1:LSC & C1:SUS restarted

Earlier this evening C1:LSC died then I hit the DAQ reload after adding an OAF channel to be recorded. No change to any model. Had to restart C1:SUS too. Reloaded burts from this morning 5am, except for C1:IOO, which I loaded from 16:07.

  5900   Tue Nov 15 22:31:39 2011 MirkoUpdateAdaptive FilteringTowards wiener filtering and improved OAFing


[Jenne, Mirko]

1. We should help the OAF by compensating for the actuator TF:

The actuator TF, from adaptive filter output to MC2, through PD, mixer, Pentek and into C1:IOO looks like this:

 TFofTheMclLoop.pdf

If we assume a white-ish error signal that the adaptive code tries to compensate for its job gets extra complicated because it has to invert this TF. So we really should compensate for that. Easiest place for that is the CORR filter directly behind the adaptive code block.

Using the TF measurement from above I used the vectfit (" /cvs/cds/caltech/apps/mDV/extra/firfit_forFotonMirkoComplex.m" ) to get fit a corresponding digital filter:

MCL_round_trip.png

If we invert swap the zeros and poles in the digital filter we get the inverted TF.
(Todo: Figure out how to invert the TF. Just switching the poles and zeros doesn't work).

2. Wiener filtering

The idea was to use the adaptive filtering only for small corrections to the wiener filtering. So we really should try to get the wiener filtering going.

Howto:

1. Get data for STS1X and GUR1X and MC_F in matlab. E.g. via ligodv
2. Check the MC was in lock the entire time.
3, Filter MC_F with the actuator TF, so the wiener filter knows about that and compensates for it
4. Calculate the wiener filter " h1winolevLigoDV.m "
5. Export the data to the workspace. It is also saved to the disc as "h1filtcoeffTS.mat". Make sure there are first the witnesses, then MC_F
6. Execute " /cvs/cds/caltech/apps/mDV/extras/LHO/firfit_for_FotonMirko.m" while one directory higher. 
7. Copy the digital filter in SOS form that is printed into the matlab command line and put it into the corresponding filter in the OAF model via foton.

With data from 11-11-15 04:00 to 05:45. Sampling freq. 256Hz. 8000 Taps => length = 30.2s. Prefiltered to notch the 60Hz line in MC_F, but not compensation the actuator TF. This results in the following wiener filter and corresponding SOS filter to be copied into foton.
STS1X:

STS1X_Wiener_filter_data_from_11-11-15.png

GUR1X:

GUR1X_Wiener_filter_data_from_11-11-15.png

Attachment 3: MCL_round_trip.fig
Attachment 6: STS1X_Wiener_filter_data_from_11-11-15.fig
Attachment 7: GUR1X_Wiener_filter_data_from_11-11-15.fig
  5899   Tue Nov 15 19:59:41 2011 SureshUpdateIOOWFS output matrix measured (open loop)

Quote:

Quote:

 The scripts used to make the WFS outmatrix measurement live in /cvs/cds/rtcds/caltech/c1/scripts/MC/WFS

 I assume you mean /opt/rtcds/caltech/c1/scripts/MC/WFS.

As I've tried to reitterate many times: we do not use /cvs/cds anymore.  Please put all new scripts into the proper location under /opt/rtcds.

 Yes the files are in /opt/rtcds/caltech/c1/scripts/MC/WFS.

I just went to wherever the 'scripts' alias takes me, found the 'pwd' and did a cp+paste of the path.   I checked to be sure that 'scripts' takes me to /opt/rtcds/caltech/c1/scripts/.

So why does the pwd show /cvs/cds.... instead of /opt/rtcds  ?

 

  5898   Tue Nov 15 16:25:38 2011 steveUpdatePSLIOO angle qpd centered

This moring I centered the IOO Angle QPD. The IOO Pos QPD was not centered. The existing layout does not allow the beam centering of the Pos qpd without misaligning the MC

input. We have to add an aditional steering mirror. I will do that tomorrow morning.

  5897   Tue Nov 15 16:16:34 2011 steveUpdatePSLenclosure interlocks are working on all sliding doors

Ben and Sam came over to fix one of  the east side  sliding  door sensor that had to be moved from the inside  to outside on the enclosure.

We turned off the 2w Innolight for 20minutes. The power is back on, the  PMC and MC locked itself immediately.

  5896   Tue Nov 15 15:56:23 2011 jamieUpdateCDSdataviewer doesn't run

Quote:

Dataviewer is not able to access to fb somehow.

I restarted daqd on fb but it didn't help.

Also the status screen is showing a blank while form in all the realtime model. Something bad is happening.

 So something very strange was happening to the framebuilder (fb).  I logged on the fb and found this being spewed to the logs once a second:

[Tue Nov 15 15:28:51 2011] going down on signal 11
sh: /bin/gcore: No such file or directory
[Tue Nov 15 15:28:51 2011] going down on signal 11
sh: /bin/gcore: No such file or directory
[Tue Nov 15 15:28:51 2011] going down on signal 11
sh: /bin/gcore: No such file or directory
[Tue Nov 15 15:28:51 2011] going down on signal 11
sh: /bin/gcore: No such file or directory
[Tue Nov 15 15:28:51 2011] going down on signal 11
sh: /bin/gcore: No such file or directory
[Tue Nov 15 15:28:51 2011] going down on signal 11
sh: /bin/gcore: No such file or directory
[Tue Nov 15 15:28:51 2011] going down on signal 11
sh: /bin/gcore: No such file or directory
[Tue Nov 15 15:28:51 2011] going down on signal 11
sh: /bin/gcore: No such file or directory
[Tue Nov 15 15:28:51 2011] going down on signal 11
sh: /bin/gcore: No such file or directory

Apparently /bin/gcore was trying to be called by some daqd subprocess or thread, and was failing since that file doesn't exist.  This apparently started at around 5:52 AM last night:

[Tue Nov 15 05:46:52 2011] main profiler warning: 1 empty blocks in the buffer
[Tue Nov 15 05:46:53 2011] main profiler warning: 0 empty blocks in the buffer
[Tue Nov 15 05:46:54 2011] main profiler warning: 0 empty blocks in the buffer
[Tue Nov 15 05:46:55 2011] main profiler warning: 0 empty blocks in the buffer
[Tue Nov 15 05:46:56 2011] main profiler warning: 0 empty blocks in the buffer
...
[Tue Nov 15 05:52:43 2011] main profiler warning: 0 empty blocks in the buffer
[Tue Nov 15 05:52:44 2011] main profiler warning: 0 empty blocks in the buffer
[Tue Nov 15 05:52:45 2011] main profiler warning: 0 empty blocks in the buffer
GPS time jumped from 1005400026 to 1005400379
[Tue Nov 15 05:52:46 2011] going down on signal 11
sh: /bin/gcore: No such file or directory
[Tue Nov 15 05:52:46 2011] going down on signal 11
sh: /bin/gcore: No such file or directory

The gcore I believe it's looking for is a debugging tool that is able to retrieve images of running processes.  I'm guessing that something caused something int the fb to eat crap, and it was stuck trying to debug itself.  I can't tell what exactly happend, though.  I'll ping the CDS guys about it.  The daqd process was continuing to run, but it was not responding to anything, which is why it could not be restarted via the normal means, and maybe why the various FB0_*_STATUS channels were seemingly dead.

I manually killed the daqd process, and monit seemed to bring up a new process with no problem.  I'll keep an eye on it.

  5895   Tue Nov 15 15:16:04 2011 kiwamuUpdateCDSdataviewer doesn't run

Dataviewer is not able to access to fb somehow.

I restarted daqd on fb but it didn't help.

Also the status screen is showing a blank white form in all the realtime model. Something bad is happening.

blank.png

JAMIEEEE !!!!

  5894   Tue Nov 15 12:25:38 2011 kiwamuUpdateGreen LockingY arm ALS : beat-note free run fluctuation

Locking activity last night :

  The free run beat-note in 532 nm has been measured.

However I couldn't close the ALS loop somehow.

Every time I tried closing the loop it broke the Y end PDH lock in a couple of minutes.

 

noise_budget.png

 

 (Things to be done)

   1.  Optimization of the Y end PDH servo loop

      1.1 Measurement of the arm fluctuation => to allow re-designing the servo shape
      1.2 Preparation of PDH box, and temporary SR560 servo
      1.3 Sanity checks on the modulation depth, reflectivity, PD dark noise and etc.,
      1.4 Make the servo more robust
      1.5 Some modifications on the medm screens
      1.6 Activation of the temperature feedback through the realtime digital control

   2. Refinement of the broadband RFPD setup

      2.1 Investigation of the peak source => there was a relatively big peak around 50 MHz or so.
      2.2 Noise characterization of the frequency detection system
      2.3 Nicer routing of some cables.
      2.4 Make two-more ADC channel connectors
      2.5 Power budget on the PSL beat-note setup => estimate the expected RF level of the beat-note
      2.6 Realignment of the PSL doubling and resetting of the doubling oven temperature
     
  3. Noise budgeting
 
     3.1 IR locked condition  => measure the noise in the green beat-note system.
     3.2 ALS engaged condition
          3.2.0 shot noise
          3.2.1 ADC noise
          3.2.2 PD dark noise
          3.2.3 freq. discriminator noise
          3.3.4 DAC noise through the coil-magnet actuators
          3.3.5 End laser suppression
          3.3.6 Intensity noise
          3.3.7 Thermo-elastic noise
          3.3.8 Thermo-refractive noise

 

  5893   Tue Nov 15 09:51:04 2011 ZachUpdateGreen LockingY end PDH lock : UGF at 17 kHz

Also the servo shape formed by Newfocus LB1005 looks too simple : we should have a more sophisticated servo filter (i.e. PDH box!!).

 As promised, I will get on this this week.

  5892   Tue Nov 15 01:44:36 2011 SureshUpdateIOOMC WFS Servo: Open loop gain

Quote:

Somehow, I generically don't like the idea of lead filters for the WFS loops. We don't really need so much bandwidth. I think you should include with the servo measurements, a servo model ( on the same plot ) that matches the loop shape.

For example, this means including the 28 Hz ELP in the MC1/3 hardware and MC2 ASCPIT/YAW digital filter banks. BY comparing the model v. measurement we can determine if the cross-coupling due to imperfect output matrix is very serious or not.

In the measurements, the loop with the most low frequency gain looks the most promising.

WFS1_PIT servo replotted with foton data overlaid:

I included the following filters in foton:

1) Integrator: zpk([0.8],[0],0.8,"n")

2) zpk([0.8],[100,100],1,"n")

3) zpk([1:10],[3,30],1,"n")

4) ELP28

I have unwound the phase by adding or subtracting 180 to portions of the phase data.

And here is the plot for WFS1_PIT.  I will repeat this process for the other three WFS loops tomorrow.

WFS1PIT_OL_gain.png

 

  5891   Tue Nov 15 00:00:15 2011 SureshUpdateIOOMC was realigned to remove beam clipping and to accommodate PZT1 range

[Kiwamu, Suresh]

The MC was realigned to readjust the input beam direction in pitch such that the clipping of the beam at the PSL table reduced and the railing of the PZT1 is avoided.

The current spot positions are given below on the last row:

 

 Date  #### MC1P MC2P MC3P MC1Y MC2Y MC3Y
03Nov2011   0.1354 -0.2522 -0.1383 -1.0893 0.7122 -1.5587
04Nov2011   4.0411 4.4994 3.5564 -1.4170 -0.2606 -1.7109
08Nov2011   4.7341 4.8794  4.3907 1.3542 -3.0508 -1.7167
10Nov2011    1   3.9944 3.7676 6.1001 -1.3058 -3.8087 -1.6418
11Nov2011    1  3.8542 3.6831 3.0418 -0.8383 0.1550 -2.3841
11Nov2011    2    3.6876 2.7429 2.7830 -1.6250 -0.0386 -1.6346
14Nov2011    1 5.9412 2.7658 5.4806 -4.7676 0.7778 2.2053

 

We have quite a lot of decentering in the MC which we must try to remove by parallel transporting the beam in Pitch and Yaw..

At the current settings we might be clipping on the Faraday Isolator as we had estimated that we can allow atmost a 2mm offset in spot positions due to this constraint.

 

 

 

  5890   Mon Nov 14 22:56:31 2011 kiwamuUpdateGreen LockingY end PDH lock : UGF at 17 kHz

[Tomotada / Kiwamu]

  The open loop transfer function of the Y end PDH loop was remeasured : the UGF was found to be at 17 kHz.

The phase margin at the UGF was about 27 deg.

YendOLTF.png

 

While the measurement we noticed that the modulation onto the laser PZT was too big

and it was creating a big AM on the reflected light with an amplitude of a few mV.

So we put a 20 dB attenuator to decrease the modulations and the reflected light became much quitter.

Also the servo shape formed by Newfocus LB1005 looks too simple : we should have a more sophisticated servo filter (i.e. PDH box!!).

  5889   Mon Nov 14 21:22:48 2011 ranaConfigurationComputersprimetime RSYNC slowing down NODUS

nodus:elog>w; who ; date
  9:20pm  up 44 day(s),  5:14,  5 users,  load average: 0.29, 1.04, 1.35
User     tty           login@  idle   JCPU   PCPU  what
controls pts/1         9:18pm            5         -tcsh
controls pts/2         2:37pm  6:39  25:02  25:02  /opt/rsync/bin/rsync -avW /cvs/c
controls pts/3         9:14pm                      w
controls pts/4         4:20pm  1:56   5:02   5:02  ssh -X rosalba
controls pts/8         8:23pm    47   4:03         -tcsh
controls   pts/1        Nov 14 21:18    (pianosa.martian)
controls   pts/2        Nov 14 14:37    (ldas-cit.ligo.caltech.edu)
controls   pts/3        Nov 14 21:14    (rosalba)
controls   pts/4        Nov 14 16:20    (192.168.113.128)
controls   pts/8        Nov 14 20:23    (gwave-103.ligo.caltech.edu)
Mon Nov 14 21:20:48 PST 2011

we will ask the man to stop running backups at this time of night...

  5888   Mon Nov 14 17:01:14 2011 kiwamuUpdateGreen LockingALS feedback on MC2

Leaving a note on the ALS feedback before I forget:

The MC2 suspension needs to have an input for the ALS feedback in the realtime model like ETMs.

  5887   Mon Nov 14 15:46:14 2011 steveUpdateVACTP2's forepump changed

The foreline pressure of TP2 was 1.4 Torr this morning. This drypump worked well for ten months.

Recently rebuilt  drypump with new seal was swapped in.

This is how you do it: close  V1,  V4 and turn off TP2. Replace drypump and start up TP2

Set pump speed to 50 K rpm and open V4 to TP1 Note that the Maglev was not turned off  because V4 was closed off only 5-10 minutes.

Open V1 the status is Vac Normal.

TP2 is rotating at  50K rpm, current pick up 0.2A,  the temp is 26C and its foreline pressure 33 mTorr

Attachment 1: drypump.png
drypump.png
  5886   Mon Nov 14 12:16:41 2011 JenneUpdateComputersOAF model died for unknown reason

I am meditating on the OAF, and had it running and calculating things.  I had the outputs disabled so I could take reference traces in DTT, but the Adapt block was calculating for MCL.  At some point, all the numbers froze, and the CPU meter had gone up to ~256ms.  Usually it's around ~70 or so for the configuration I had (2 witness sensors and one degree of freedom enabled....no c-code calculations on any other signals).  The "alive" heartbeat was also frozen.

I ssh'ed into c1lsc, ran ./startc1oaf in the scripts directory, and it came back just fine.

Anyhow, I don't know why it got funny, but I wanted to record the event for posterity.  I'm back to OAFing now.

  5885   Mon Nov 14 11:32:02 2011 kiwamuSummaryGeneralGoal this week

Goal of this week : Noise budgeting on the Y arm ALS

Minimum success : bring the Y arm to the resonance by using ALS  NOISE BUDGETING!!!

 => as a preparation the incident beam pointing needs to be fixed by steering the MC suspensions.

Quote from #5850http://nodus.ligo.caltech.edu:8080/40m/5850

Goal of this week :  ALS on the Y arm (DONE)

  5884   Sat Nov 12 08:09:47 2011 ranaUpdateIOOMC WFS Servo: Open loop gain

Somehow, I generically don't like the idea of lead filters for the WFS loops. We don't really need so much bandwidth. I think you should include with the servo measurements, a servo model ( on the same plot ) that matches the loop shape.

For example, this means including the 28 Hz ELP in the MC1/3 hardware and MC2 ASCPIT/YAW digital filter banks. BY comparing the model v. measurement we can determine if the cross-coupling due to imperfect output matrix is very serious or not.

In the measurements, the loop with the most low frequency gain looks the most promising.

  5883   Sat Nov 12 03:46:55 2011 SureshUpdateIOOMC WFS Servo: Open loop gain

[Mirko, Suresh]

I closed the WFS loops and measured the transfer function from IN2 to IN1 testpoints on the WFS1_PIT filterbank. 

We looked at the filter shape consisting of

1) Integrator: zpk([0.8],[0],0.8,"n")

2) zpk([0.8],[100,100],1,"n")

3) zpk([1:10],[3,30],1,"n")

The combined filter shape (along with an added pendulum filter, zpk([ ],0.8,1,"n")  ) is given below

WFS1_PIT_servo_filtershape_20111111_1.png

 

The OL Transfer function measured for WFS1_PIT loop is

WFS1_PIT_servo_OLG_20111111_1.png

 The blue reference is a measurement  without the third "45 deg" filter in the list above.  Without it the UGF is around 1.5Hz and increasing the gain results in additional noise from the servo bump seen in the earlier elog .  With it the UGF is around 3Hz.

The supression of the error signal is shown here

 Error_signal_WFS1_PIT_20111111_1.png

The other WFS loops are expected to have a similar behaviour with the exception of the MC2 QPD channels.  I will measure their OLTF shortly and then proceed with the inclusion of the QPD sensors into the WFS system.

 

 

  5882   Sat Nov 12 02:46:13 2011 DenUpdateAdaptive Filteringstacks and ground

We measured the coherence between the seismometer near the MC2 stack and accelerometers on the vacuum tank where MC2 is. Because accelerometers produce small signals at low frequencies, which are comparable with adc noise, we  amplified the accelerometer signal by a factor of 20. We could not do it more because though adc has 40 V range, the black box that follows the channel sockets can transmit only 2.5 V max amplitude signal. Probably, this was done because old adc accepted 2 V max amplitude.

ground_stack.jpg

ground_stack_coherence.jpg


We were able to found some coherence at 0.1-1 Hz though the accelerometer signal is rather noisy. So to consider stack as a noise source is still possible. This measurement should be better done with two seismometers, one on the floor, the other on the stack. From the figure we can also see that tilt affects the x and y seismometer signals from 0.1 Hz. Green line (z-component) is much lower that red and blue lines (x and y). Tilt affects on horisontal axes of the seismometer much more than on vertical.

What we also think about is that at low frequencies mirrors start to move approximately the same and seismometers can help us to figure out small reletive displacement of the mirrors which form the MC length. We can estimate the critical frequency by presenting the ground motion as interference of surface waves with different velocities and amplitudes. For only 1 wave we have for the relation of MC length to the seismometer read out  ~sin (2*pi*f/v*L). f - frequency of the wave, v -speed, L - length between the mirrors. We can see that below 1 Hz we have ~sin (f/2). At this point seismometer signal could lose coherence with MC length signal. We could try to subtract seismometer signals from corresponding axes, but gur1 and sts1 has different calibrations. Moreover, the noise floor of the seismometers might not allow us to measure the differential signal. We'll try to simulate this scenario and find seismometer calibration or measure it. We are basicly interested only in the ratio of calibraion fucntions of 2 seismometers.

  5881   Sat Nov 12 02:44:18 2011 SureshUpdateComputersC1IOO front end suddenly froze. Was restarted remotely

Quote:

[Koji Suresh]

No one was messing with the c1ioo or any other machine.   The medm screens for WFS and MC alignment froze while I was working on Rossa.

There were number of red lights pertaining to c1ioo machine on the CDS_FR_STATUS screen.  So we logged into c1ioo   from Rossa and restarted it with 'sudo shutdown -r now'.  It came back up but the C1IOO_MC_TRANS_SUM, P and Y signals were not available on the C1IOO LOCKMC screen.

I saw several messages similar to the one here


Sat Nov 12 02:09:14 PST 2011

  medmCAExceptionHandlerCb: Channel Access Exception:
  Channel Name: Unavailable
  Native Type: Unavailable
  Native Count: 0
  Access: Unavailable
  IOC: Unavailable
  Message: Virtual circuit disconnect
  Context: c1ioo.martian:43553
  Requested Type: TYPENOTCONN
  Requested Count: 0
  Source File: ../cac.cpp
  Line number: 1126

 

The MC autolocker script wasnt running.  The heartbeat bit was not blinking on the MC_LOCKMC screen.  So we manually restarted the script.  Hopefully it will return to normal operation.

I restarted the fb at Sat Nov 12 02:12:19 PST 2011  in an attempt to see this resolves the problem.

It didnt.

 

 The problem was resolved after I burtrestored (c1mcs c1ioo and c1rfm) epics snapshots.

 

  5880   Sat Nov 12 02:27:00 2011 SureshUpdateComputersC1IOO front end suddenly froze. Was restarted remotely

[Koji Suresh]

No one was messing with the c1ioo or any other machine.   The medm screens for WFS and MC alignment froze while I was working on Rossa.

There were number of red lights pertaining to c1ioo machine on the CDS_FR_STATUS screen.  So we logged into c1ioo   from Rossa and restarted it with 'sudo shutdown -r now'.  It came back up but the C1IOO_MC_TRANS_SUM, P and Y signals were not available on the C1IOO LOCKMC screen.

I saw several messages similar to the one here


Sat Nov 12 02:09:14 PST 2011

  medmCAExceptionHandlerCb: Channel Access Exception:
  Channel Name: Unavailable
  Native Type: Unavailable
  Native Count: 0
  Access: Unavailable
  IOC: Unavailable
  Message: Virtual circuit disconnect
  Context: c1ioo.martian:43553
  Requested Type: TYPENOTCONN
  Requested Count: 0
  Source File: ../cac.cpp
  Line number: 1126

 

The MC autolocker script wasnt running.  The heartbeat bit was not blinking on the MC_LOCKMC screen.  So we manually restarted the script.  Hopefully it will return to normal operation.

I restarted the fb at Sat Nov 12 02:12:19 PST 2011  in an attempt to see this resolves the problem.

It didnt.

 

  5879   Sat Nov 12 02:00:36 2011 MirkoUpdateAdaptive FilteringMC-F and other signals

Regarding http://nodus.ligo.caltech.edu:8080/40m/5867 and http://nodus.ligo.caltech.edu:8080/40m/5869 :

MC_F signal:

The measurements on p. 5867 were done using the ADC attached to the PEM computer. There was a big difference between the MC_F signal recorded directly after the server board and the signal just before the FSS board as recorded by a PEM channel.
To understand how this happens we measured the signal at different places with a spec. analyzer:

1. WIth a locked MC measuring the signal just before the PEM ADCs (meaning after a 60ft BNC cable)
2. Same position, but unlocked and seemingly dark MC
3. Locked MC, signal just before the FSS box
4. MC_F signal that is usually going into the Pentek Generic board and is recorded in C1:IOO

Compare_signals_at_all_places.png

=> The 60ft BNC cable adds a considerable amount of noise, but doesn't fundamentally change the signal. It is weird that the signal is white from approx. 4Hz on.
Due to Jenne's measurement ( http://nodus.ligo.caltech.edu:8080/40m/5848 ) we know the TF from MCL through PD, mixer Pentek and into C1:IOO looks like this:
OAF-MCL-Delay-9Nov2011.pdf

This is with the double HP from 15Hz on that should be in the Pentek. So one might expect a less white signal going into the FSS board...

 PEM ADCs

The dark noise in the PEM ADCs is actually a factor 10 higher than in the IOO ADCs. Still ok wrt the the seismometers.
We also tried to measure essentially the dark noise of the whole seismometer readout (seismometer box, then ADC). That seemed ok, but is of limited value since the seismometer electronics behave a bit strange when there is no seismometer connected.

Channels_attached_to_the_PEM_ADC.png

Attachment 3: Compare_signals_at_all_places.fig
Attachment 5: Channels_attached_to_the_PEM_ADC.fig
  5878   Fri Nov 11 22:07:43 2011 SureshUpdateIOOTried to recover the MC alignment of 4th Nov: partial success, PSL beam clipping

I have recovered the yaw values pretty much .  As the PZT1 rails in this direction perhaps this is the more relevant of the two alignments.  The beam is translated in the vertical direction, but this can be easily corrected by changing the pitch of MC2

However note that if the WFS are switched on .. MC is going to follow the PSL beam. 

 

 

 Date  #### MC1P MC2P MC3P MC1Y MC2Y MC3Y
03Nov2011   0.1354 -0.2522 -0.1383 -1.0893 0.7122 -1.5587
04Nov2011   4.0411 4.4994 3.5564 -1.4170 -0.2606 -1.7109
08Nov2011   4.7341 4.8794  4.3907 1.3542 -3.0508 -1.7167
10Nov2011    1   3.9944 3.7676 6.1001 -1.3058 -3.8087 -1.6418
11Nov2011    1  3.8542 3.6831 3.0418 -0.8383 0.1550 -2.3841
11Nov2011    2    3.6876 2.7429 2.7830 -1.6250 -0.0386 -1.6346

 

 

  5877   Fri Nov 11 21:09:30 2011 SureshUpdateSUSMC2 is being a little wild...WFS to blame

Quote:

Mirko and  Den are measuring MC_F, which they will report about later, but I noticed that MC2 is totally crazy right now.  It shouldn't matter that they are doing things (like unplugging the feedback to the PSL's PZT), because we actuate on the laser, not on the MC.  I disabled the MC autolocker before they started working. 

Anyhow, somehow MC2 got kicked up (whatever, that happens), but it won't re-damp.  I think it's the WFS.  The yaw output from the WFS is truely crazy. 

I have disabled the WFS output / ASC input on the MC SUS screens, and MC2 was then able to damp.  My disabling only the MC2 WFS input at first kicked up MC1 and 3, so I disabled all of the WFS stuff, and all 3 MC mirrors are again happy. 

SURESH: FIX ME!  (signed, The WFS)

WFSscreenshot.png

 

The Problem

Turning off the WFS servo loops usually should be done using the 'mcwfsoff' script.  The script takes care of switching off the integrators and Clears the History.  

'mcdown' and 'mcup' scripts run the 'mcwfsoff' and 'mcwfson' scripts so when the MC unlocks the WFS servos are shutdown and restarted properly.  However if the MC autolocker script is suspended by pressing the Enable/Disable switch in the LOCKMC screen and then the MC unlocks, it results in the WFS servo integrators accumulating large values.  If these values are passed through the ASC filter banks the optic will get a pretty huge kick.

The Solution

I have added some indicators which will let us know if the WFS Servo Filter outputs are larger than +/-1000.  When engaging the WFS loops the user has to take care to Clear History in the servo filter banks if these indicators are steadily Red.   before engaging the WFS Servo loops ensure that the servo filter outputs are zeroes. 

Koji and I discussed whether it would be useful to run the 'mcwfsoff' script when the Disable button is pressed in the autolocker.  His recommendation is that we should keep  the autolocker script simple and that user has to be cautious when switching on the WFS servos and when directing the ASC outputs to the suspensions.

LOCKMC_screen.png

 

  5875   Fri Nov 11 14:55:47 2011 kiwamuSummaryLSCCharacterization of the Power Recycled Michelson : take 2

Quote from #5851

The recycling gain is determined by the optical configuration and the optical loss in the cavity.

How much is the actual recycling gain? And how does it affect the signal extraction?

 As Koji pointed out I made a wrong definition on the recycling gain of PRMI (Power-Recycled Michelson Interferomter).

In the correct definition the estimated recycling gain is 15.
In order to answer Koji's second question,which is about the effect on the signal extraction,
I need to scratch my head for a while.
( Give me some time..)
 
The value what I called "Recycling gain" must have been called "measured power build up" or something like that.
For clarity I put the definitions of the quantities.
    Recycling gain :      rec_gain.png 

   Reflectivity of PRMI (measured by REFLDC): refl.png

    Power build up (measured by POY DC) : pbu.png

    Mode Matching (MM) efficiency :  MM.png

    Loss in the PRMI cavity : loss.png

 


 (Results of Measurement and Estimation)

     Estimated recycling gain = 15

     Estimated MM efficiency = 47.4%

     Estimated Loss = 5.3%

     Measured power build Up = 7

     Measured reflectivity of PRMI = 0.5

  5874   Fri Nov 11 13:35:19 2011 KatrinUpdateGreen LockingFeedback to ETMY

[Kiwamu, Katrin]

20111111_freq_supp.png

 

Red and blue curves: frequency fluctuation of the beat node between PSL and YARM laser.

Green and broen curves: Actuation on ETMY.  In ALS_CONTROL.adl  ETMY filter bank 4 and 5 were switched on. Gain was 0.3

 

Nice reduction of the frequency fluctuation.

 

Y axis is in volts^2 per counts. In order to go to MHz/sqrt(Hz) you have to take the square root and then times [20Volts/(2^16)counts]*[10Hz/0.04V].

 

Started to scan the cavity, but this didn't work. Green light all out of lock. IR beam was badly aligned to cavity. Now, my time is over and I have to leave you.

Thanks, for your help and the nice time.

ELOG V3.1.3-